# Optimizing Parameters of the LinDistFlow Power Flow Approximation for Distribution Systems

### v3

@@ Abstract
Index Terms:
DistFlow, LinDistFlow, machine learning, parameter optimization, distribution systems, hosting capacity.
@@ I Introduction
Power flow models relating power injections, line flows, and voltages are central to power system design and operation [1]. The AC power flow equations accurately model these relationships. Incorporating the AC power flow equations into optimization problems introduces significant computational challenges due to these equations’ nonlinearity, even for the radial networks of typical distribution systems [2]. These challenges are particularly relevant to problems that are run online to inform real-time decisions [3, 4], consider uncertainties [5], and model discrete decisions [6, 7, 8, 9]. To address these challenges, engineers frequently use power flow approximations that trade accuracy for tractability [10].
@@
Building on ideas from recently developed “adaptive” power flow approximations [28, 29, 30, 31, 32, 33, 34], and more recent advancements focusing on areas like second-order sensitivity insights [35] and sample-based piecewise linear approaches [36], this paper introduces a parameter optimization algorithm for the LinDistFlow model (and its three-phase extension, LinDist3Flow) to significantly enhance its accuracy. Our method optimizes key coefficient parameters, specifically the elements of the diagonal matrices $D_{r}$ and $D_{x}$ which represent effective line resistances and reactances, and newly introduced bias parameters. These bias parameters are vector offsets $\boldsymbol{\rho}$ for active power injections ($p$), $\boldsymbol{\varrho}$ for reactive power injections ($q$), and $\boldsymbol{\gamma}$ for squared voltage magnitudes ($v$). These parameters, which form the basis of our Optimized LinDistFlow (OLDF) formulation detailed further in Section III, are tuned via an offline training process. This allows the OLDF model to better emulate the nonlinear DistFlow model, thereby more accurately capturing the net impact of complex physical effects, such as those related to power losses, which are simplified in traditional LinDistFlow formulations.
@@
  • A systematic framework for optimizing LinDistFlow and LinDist3Flow to significantly enhance accuracy against nonlinear DistFlow by tuning coefficient and bias parameters using sensitivity analysis and the TNC method.
@@
  • Demonstration of the optimized LinDistFlow’s efficacy in practical applications, yielding more reliable and feasible solutions than traditional LDF in complex engineering tasks like hosting capacity analysis.
@@ II Power Flow Modeling
[…]
II-A DistFlow Model for Single-Phase Equivalent Networks
@@
$v_{n}=|V_{n}|^{2},\quad\ell_{n}=|I_{n}|^{2},$ $\forall n\in\mathcal{N},$ (1a)
$\sum_{k:n\rightarrow k}P_{k}=p_{n}+P_{n}-r_{n}\ell_{n},$ $\forall n\in\mathcal{N},$ (1b)
$\sum_{k:n\rightarrow k}Q_{k}=q_{n}+Q_{n}-x_{n}\ell_{n},$ $\forall n\in\mathcal{N},$ (1c)
$v_{n}=v_{\pi_{n}}-2(r_{n}P_{n}+x_{n}Q_{n})+(r_{n}^{2}+x_{n}^{2})\ell_{n},$ $\forall n\in\mathcal{N},$ (1d)
$v_{\pi_{n}}\ell_{n}=P_{n}^{2}+Q_{n}^{2},$ $\forall n\in\mathcal{N}.$ (1e)
Equation (1a) defines variables for the squared voltage magnitudes $v_{n}$ and squared current flow magnitudes $\ell_{n}$ which are used in place of the voltage phasors $V_{n}$ and current phasors $I_{n}$. Equations (1b) and (1c) correspond to active and reactive power balance at each bus. Equation (1d) models the change in squared voltage magnitudes across lines. Equation (1e) uses the definition of apparent power to relate the squared voltages $v_{n}$, squared currents $\ell_{n}$, and squared apparent power $P_{n}^{2}+Q_{n}^{2}$. The DistFlow model is nonlinear due to (1e).
@@
II-B Traditional LinDistFlow Approximation
@@
$\sum_{k:n\rightarrow k}P_{k}\approx p_{n}+P_{n},$ $\forall n\in\mathcal{N},$ (2a)
$\sum_{k:n\rightarrow k}Q_{k}\approx q_{n}+Q_{n},$ $\forall n\in\mathcal{N},$ (2b)
$v_{n}\approx v_{\pi_{n}}-2(r_{n}P_{n}+x_{n}Q_{n}),$ $\forall n\in\mathcal{N}.$ (2c)
@@ III Optimized LinDistFlow Approximation (OLDF)
Appropriate selection of parameter values for $\mathbf{D}_{r}$, $\mathbf{D}_{x}$, $\boldsymbol{\gamma}$, $\boldsymbol{\rho}$, and $\boldsymbol{\varrho}$ can significantly improve the LinDistFlow approximation’s accuracy. By adjusting the effective nodal power injections $(\mathbf{p}+\boldsymbol{\rho})$ and $(\mathbf{q}+\boldsymbol{\varrho})$, the parameters $\boldsymbol{\rho}$ and $\boldsymbol{\varrho}$ allow the linear model to implicitly account for the voltage impact of system characteristics not directly captured in the basic LinDistFlow equations, such as line losses. We next introduce a machine learning-inspired algorithm which optimizes these parameters to reduce the discrepancy between the voltages output by the LinDistFlow approximation (III) and the DistFlow model (1).
@@ III-A Formulation of the Loss Function
We define a loss function, $\mathcal{L}$, as the mean square error between the voltage solutions of the DistFlow model ($\mathbf{v}_{m}^{DF}$) and our optimized LinDistFlow approximation ($\mathbf{v}_{m}^{OLDF}$) across a given set of load scenarios =$\mathcal{M}=\{1,2,\ldots,S\}$:
$\mathcal{L}(\mathbf{D}_{r},\mathbf{D}_{x},\boldsymbol{\gamma},\boldsymbol{\rho},\boldsymbol{\varrho})=\frac{1}{|\mathcal{M}||\mathcal{N}^{\prime}|}\sum_{m\in\mathcal{M}}||\mathbf{v}_{m}^{OLDF}-\mathbf{v}_{m}^{DF}||^{2}_{2},$
@@
$\min_{\mathbf{D}_{r},\mathbf{D}_{x},\boldsymbol{\gamma},\boldsymbol{\rho},\boldsymbol{\varrho}}\mathcal{L}(\mathbf{D}_{r},\mathbf{D}_{x},\boldsymbol{\gamma},\boldsymbol{\rho},\boldsymbol{\varrho}).$ (6)
@@ III-B Parameter Sensitivity Analysis
$\mathbf{g}^{\eta}=\frac{2}{|\mathcal{M}||\mathcal{N}^{\prime}|}\sum_{m\in\mathcal{M}}\left.\frac{\partial\mathbf{v}^{OLDF}}{\partial\eta}\right|_{\mathbf{v}^{OLDF}_{m}}\Big{(}\mathbf{v}_{m}^{OLDF}-\mathbf{v}_{m}^{DF}\Big{)},$ (7a)
@@
$\frac{\partial\mathbf{v}^{OLDF}}{\partial\mathbf{D}_{r}}$ $=2\Big{(}\mathbf{A}^{-\top}(\mathbf{p}+\boldsymbol{\rho})\Big{)}^{\top}\mathbf{A}^{-1},$ (7b)
$\frac{\partial\mathbf{v}^{OLDF}}{\partial\mathbf{D}_{x}}$ $=2\Big{(}\mathbf{A}^{-\top}(\mathbf{q}+\boldsymbol{\varrho})\Big{)}^{\top}\mathbf{A}^{-1},$ (7c)
@@ III-C Implementation of the Optimization Solution
The gradients in Section III-B enable the application of gradient-based optimization methods such as TNC [45, 46] to solve the parameter optimization problem (6). The choice of TNC as opposed to other gradient-based optimization methods such as Broyden-Fletcher-Goldfarb-Shanno (BFGS) and limited-memory BFGS (L-BFGS[45, 46] is based on the TNC method’s superior scalability in our empirical testing. Specifically, for our parameter optimization problem, TNC demonstrated superior practical performance over L-BFGS variants, converging with less computational time, particularly for larger test cases. This advantage in handling large-scale systems is consistent with our findings in related work [37], where TNC also proved more effective.
@@
           $\nabla\mathcal{L}(\mathbf{x}_{k})$$=\mathbf{g}=[{\mathbf{g}^{D_{r}}}^{\top},{\mathbf{g}^{D_{x}}}^{\top},{\mathbf{g}^{\gamma}}^{\top},{\mathbf{g}^{\rho}}^{\top},{\mathbf{g}^{\varrho}}^{\top}]^{\top}$
@@
           $\alpha_{2}$: Curvature condition constant, between $\alpha_{1}$ and $1$
@@ 6 $k\leftarrow 0$
[…]
@@ IV Three-Phase LinDistFlow (LinDist3Flow)
where $\mathbb{V}=\begin{bmatrix}v^{a}_{1}&v^{b}_{1}&v^{c}_{1}&\cdots&v^{a}_{n}&v^{b}_{n}&v^{c}_{n}\end{bmatrix}^{\top}$ is the vector of squared of voltage magnitudes for each phase $(a,b,c)$, $\mathbb{A}_{3}$ is the network incidence matrix, $\text{bdiag}(\,\cdot\,)$ is the block diagonal operator, $\mathbb{P}=\begin{bmatrix}p^{a}_{1}&p^{b}_{1}&p^{c}_{1}&\cdots&p^{a}_{n}&p^{b}_{n}&p^{c}_{n}\end{bmatrix}^{\top}$ and $\mathbb{Q}=\begin{bmatrix}q^{a}_{1}&q^{b}_{1}&q^{c}_{1}&\cdots&q^{a}_{n}&q^{b}_{n}&q^{c}_{n}\end{bmatrix}^{\top}$ are the active and reactive power injection vectors, and the $\mathbb{H}$ matrices for each line $(i,j)\in\mathcal{E}$ are:
@@
$\mathcal{L}(\mathbb{H}^{P},\mathbb{H}^{Q},\boldsymbol{\gamma}_{3},\boldsymbol{\rho}_{3},\boldsymbol{\varrho}_{3})=\frac{1}{|\mathcal{M}||3\mathcal{N}^{\prime}|}\sum_{m\in\mathcal{M}}||\mathbb{V}_{m}^{OLDF}-\mathbb{V}_{m}^{DF}||^{2}_{2},$
@@
$\mathbf{g}^{\eta}=\frac{2}{|\mathcal{M}||3\mathcal{N}^{\prime}|}\sum_{m\in\mathcal{M}}\left.\frac{\partial\mathbb{V}^{OLDF}}{\partial\eta}\right|_{\mathbb{V}^{OLDF}_{m}}\Big{(}\mathbb{V}_{m}^{OLDF}-\mathbb{V}_{m}^{DF}\Big{)},$ (12a)

### v4

@@ I Introduction
Power flow models relating power injections, line flows, and voltages are central to power system design and operation [1]. While all mathematical models of physical systems are inherently approximations, the AC power flow equations are the established, high-fidelity benchmark used to accurately model these relationships. Incorporating the AC power flow equations into optimization problems introduces significant computational challenges due to these equations’ nonlinearity, even for the radial networks of typical distribution systems [2]. These challenges are particularly relevant to problems that are run online to inform real-time decisions [3, 4], consider uncertainties [5], and model discrete decisions  [6, 7, 8, 9]. To address these challenges, engineers frequently use power flow approximations that trade accuracy for tractability [10].
@@
Building on ideas from recently developed “adaptive” power flow approximations [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], and related advancements focusing on areas like second-order sensitivity insights [43] and sample-based piecewise linear approaches [44, 45], this paper introduces a parameter optimization algorithm for the LinDistFlow model (and its three-phase extension, LinDist3Flow) to significantly enhance its accuracy. Our method optimizes key coefficient parameters, specifically the elements of the diagonal matrices $D_{r}$ and $D_{x}$ which represent effective line resistances and reactances, and newly introduced bias parameters. These bias parameters are vector offsets $\boldsymbol{\rho}$ for active power injections ($p$), $\boldsymbol{\varrho}$ for reactive power injections ($q$), and $\boldsymbol{\gamma}$ for squared voltage magnitudes ($v$). These parameters, which form the basis of our Optimized LinDistFlow (OLDF) formulation detailed further in Section III, are tuned via an offline training process. This allows the OLDF model to better emulate the nonlinear DistFlow model, thereby more accurately capturing the net impact of complex physical effects, such as those related to power losses, which are simplified in traditional LinDistFlow formulations.
@@
  • A systematic framework for optimizing LinDistFlow and LinDist3Flow to significantly enhance accuracy to better match nonlinear DistFlow by tuning coefficient and bias parameters using sensitivity analysis and the TNC method.
@@
  • Demonstration of the proposed OLDF’s efficacy in practical applications, yielding more reliable and feasible solutions than traditional LDF in complex engineering tasks like hosting capacity analysis.
@@ II Power Flow Modeling
II-A Notation
We first establish notation. Let operators $(\,\cdot\,)^{-1}$, $|\,\cdot\,|$, $(\,\cdot\,)^{\top}$, and $(\,\cdot\,)^{-\top}$ denote the square matrix inverse, the absolute value of a number, the transpose of a matrix/vector, and the transpose of a square matrix inverse, respectively. Let $\mathcal{N}:=\{0,1,\ldots,n\}$ and $\mathcal{E}:=\{1,\ldots,n\}$ denote the sets of buses and lines, respectively, in a distribution network, where $|\mathcal{E}|=|\mathcal{N}|-1$ for radial networks. Each bus $n\in\mathcal{N}$ has a voltage $V_{n}$. The set of all non-substation buses is denoted by $\mathcal{N}^{\prime}:=\{1,2,...,n\}$. Each bus $n\in\mathcal{N}^{\prime}$ has a parent (upstream) bus denoted as $\pi_{n}$. Each line $(\pi_{n},n)\in\mathcal{E}$ (connection between bus $n$ and its parent $\pi_{n}$) has impedance $z_{n}=r_{n}+\mathrm{j}x_{n}$ with resistance $r_{n}$ and reactance $x_{n}$, where $\mathrm{j}:=\sqrt{-1}$ (note that lines are identified by their child buses to simplify notation; see Fig. 1). Let $I_{n}$ and $S_{n}=P_{n}+\mathrm{j}Q_{n}$ denote the current and complex power flows, respectively, on line $(\pi_{n},n)$. The net power injection at bus $n$ is $s_{n}=p_{n}+jq_{n}$. Additional variables include squared voltage and current magnitudes $v_{n}=|V_{n}|^{2}$ and $\ell_{n}=|I_{n}|^{2}$. $\mathbf{D}_{r}=\text{diag}(\mathbf{r})$ and $\mathbf{D}_{x}=\text{diag}(\mathbf{x})$ are diagonal matrices of resistances and reactances, and $\mathbf{p}=[p_{1},\ldots,p_{n}]^{\top}$, $\mathbf{q}=[q_{1},\ldots,q_{n}]^{\top}$, $\mathbf{P}=[P_{1},\ldots,P_{n}]^{\top}$, and $\mathbf{Q}=[Q_{1},\ldots,Q_{n}]^{\top}$ are vectors of active/reactive power injections and flows. Define $\hat{\mathbf{A}}=[\mathbf{a}_{0}~\mathbf{A}]$ as the $|\mathcal{E}|\times|\mathcal{N}|$ branch-bus incidence matrix describing the connections between the system’s buses and branches, where $\mathbf{a}_{0}$ is the length-$|\mathcal{E}|$ vector associated with the substation bus and $\mathbf{A}$ is the reduced branch-bus matrix for all buses other than the substation bus. Let $\mathbf{v}:=[|V_{1}|^{2},\ldots,|V_{n}|^{2}]^{\top}$ represent squared voltage magnitudes and $v_{0}=|V_{0}|^{2}$ correspond to the substation bus. To enhance voltage approximation accuracy, our algorithm also optimizes a set of bias parameters. These parameters are formulated as vectors of length $|\mathcal{N}^{\prime}|$ and introduce corrective offsets to the model: $\gamma$ is an offset for the squared voltage magnitudes ($v$), $\rho$ is an offset for the active power injections ($p$), and $\varrho$ is an offset for the reactive power injections ($q$).
II-B DistFlow Model for Single-Phase Equivalent Networks
@@
$v_{n}=|V_{n}|^{2},\quad\forall n\in\mathcal{N};\quad\ell_{n}=|I_{n}|^{2},$ $\forall n\in\mathcal{N}^{\prime},$ (1a)
$\sum_{k:n\rightarrow k}P_{k}=p_{n}+P_{n}-r_{n}\ell_{n},$ $\forall n\in\mathcal{N}^{\prime},$ (1b)
$\sum_{k:n\rightarrow k}Q_{k}=q_{n}+Q_{n}-x_{n}\ell_{n},$ $\forall n\in\mathcal{N}^{\prime},$ (1c)
$v_{n}=v_{\pi_{n}}-2(r_{n}P_{n}+x_{n}Q_{n})+(r_{n}^{2}+x_{n}^{2})\ell_{n},\hskip-9.90005pt$ $\forall n\in\mathcal{N}^{\prime},$ (1d)
$v_{\pi_{n}}\ell_{n}=P_{n}^{2}+Q_{n}^{2},$ $\forall n\in\mathcal{N}^{\prime}.$ (1e)
Equation (1a) defines variables for the squared voltage magnitudes $v_{n}$ and squared current flow magnitudes $\ell_{n}$, which are used in place of the voltage phasors $V_{n}$ and current phasors $I_{n}$. Equations (1b) and (1c) correspond to active and reactive power balance at each bus. Equation (1d) models the change in squared voltage magnitudes across lines. Equation (1e) uses the definition of apparent power to relate the squared voltages $v_{n}$, squared currents $\ell_{n}$, and squared apparent power $P_{n}^{2}+Q_{n}^{2}$. The DistFlow model is nonlinear due to (1e). Note that the DistFlow model (1) is equivalent to bus-injection AC power flow models for radial systems [11, 12, 13, 47, 48].
@@
II-C Traditional LinDistFlow Approximation
@@
$\sum_{k:n\rightarrow k}P_{k}\approx p_{n}+P_{n},$ $\forall n\in\mathcal{N}^{\prime},$ (2a)
$\sum_{k:n\rightarrow k}Q_{k}\approx q_{n}+Q_{n},$ $\forall n\in\mathcal{N}^{\prime},$ (2b)
$v_{n}\approx v_{\pi_{n}}-2(r_{n}P_{n}+x_{n}Q_{n}),$ $\forall n\in\mathcal{N}^{\prime}.$ (2c)
@@ III Optimized LinDistFlow Approximation (OLDF)
While the traditional LinDistFlow model derives the matrices $D_{r}$ and $D_{x}$ directly from the physical line parameters, our approach treats them as optimizable coefficients. The justification for this is that the LinDistFlow model is an approximation of the true nonlinear physics, primarily by neglecting the impacts of line losses. Since the model’s structure is simplified, using the exact physical parameters does not guarantee the most accurate linear approximation. Instead, by optimally selecting values for the $\mathbf{D}_{r}$ and $\mathbf{D}_{x}$ coefficients, we find the “effective” resistance and reactance values that cause the simplified linear model to best match the results of the full nonlinear model over a range of operating conditions. We note that conceptually similar approaches for selecting parameter values in power flow approximations are employed in, e.g., [28, 29, 30, 31, 34, 32, 19, 36, 37, 38, 35, 33, 43, 39, 42, 45, 40, 41, 44]. In Section V, we numerically compare our optimized LinDistFlow formulation to the relevant alternative LinDistFlow approximations to show that optimizing the parameter values is an effective method for improving LinDistFlow approximation accuracy.
@@
Appropriate selection of parameter values for $\mathbf{D}_{r}$, $\mathbf{D}_{x}$, $\boldsymbol{\gamma}$, $\boldsymbol{\rho}$, and $\boldsymbol{\varrho}$ can significantly improve the LinDistFlow approximation’s accuracy. By adjusting the effective nodal power injections $(\mathbf{p}+\boldsymbol{\rho})$ and $(\mathbf{q}+\boldsymbol{\varrho})$, the parameters $\boldsymbol{\rho}$ and $\boldsymbol{\varrho}$ allow the linear model to implicitly account for the voltage impact of system characteristics not directly captured in the basic LinDistFlow equations, such as line losses. [^1] Similar adjustment parameters are often employed to improve the accuracy of DC power flow approximations [49]. We next introduce a machine learning-inspired algorithm which optimizes these parameters to reduce the discrepancy between the voltages output by the LinDistFlow approximation (III) and the DistFlow model (1).
@@ III-A Formulation of the Loss Function
We define a loss function, $\mathcal{L}$, as the mean square error between the voltage solutions of the DistFlow model ($\mathbf{v}_{m}^{DF}$) and our optimized LinDistFlow approximation ($\mathbf{v}_{m}^{OLDF}$) across a given set of load scenarios $\mathcal{M}=\{1,2,\ldots,S\}$:
[…]
@@
$\operatorname*{minimize}_{\mathbf{D}_{r},\mathbf{D}_{x},\boldsymbol{\gamma},\boldsymbol{\rho},\boldsymbol{\varrho}}\mathcal{L}(\mathbf{D}_{r},\mathbf{D}_{x},\boldsymbol{\gamma},\boldsymbol{\rho},\boldsymbol{\varrho}).$ (6)
@@ III-B Parameter Sensitivity Analysis
$\mathbf{g}^{\eta}=\frac{2}{|\mathcal{M}||\mathcal{N}^{\prime}|}\sum_{m\in\mathcal{M}}\Big(\left.\frac{\partial\mathbf{v}^{OLDF}}{\partial\eta}\right|_{\mathbf{v}^{OLDF}_{m}}\Big)^{\top}\Big(\mathbf{v}_{m}^{OLDF}-\mathbf{v}_{m}^{DF}\Big),$ (7a)
@@
$\frac{\partial\mathbf{v}^{OLDF}}{\partial\mathbf{D}_{r}}$ $=2\mathbf{A}^{-1}\text{diag}\Big(\mathbf{A}^{-\top}(\mathbf{p}+\boldsymbol{\rho})\Big),$ (7b)
$\frac{\partial\mathbf{v}^{OLDF}}{\partial\mathbf{D}_{x}}$ $=2\mathbf{A}^{-1}\text{diag}\Big(\mathbf{A}^{-\top}(\mathbf{q}+\boldsymbol{\varrho})\Big),$ (7c)
@@ III-C Implementation of the Optimization Solution
The gradients in Section III-B enable the application of gradient-based optimization methods such as TNC [50, 51] to solve the parameter optimization problem (6). The choice of TNC as opposed to other gradient-based optimization methods such as Broyden-Fletcher-Goldfarb-Shanno (BFGS) and limited-memory BFGS (L-BFGS) [50, 51] is based on the TNC method’s superior scalability in our empirical testing. Specifically, for our parameter optimization problem, TNC demonstrated superior practical performance over L-BFGS variants, converging with less computational time, particularly for larger test cases. This advantage in handling large-scale systems is consistent with our findings in related work [35], where TNC also proved more effective.
@@
    $\nabla\mathcal{L}(\mathbf{x}_{k})$$\mathbf{g}=[{\mathbf{g}^{D_{r}}}^{\top},{\mathbf{g}^{D_{x}}}^{\top},{\mathbf{g}^{\gamma}}^{\top},{\mathbf{g}^{\rho}}^{\top},{\mathbf{g}^{\varrho}}^{\top}]^{\top}$
@@
    $\alpha_{2}$: Curvature condition constant, between $\alpha_{1}$ and 11
@@ 6 $k\leftarrow 0$
8while $k\leq\textit{max\_iter}$ and $\|\nabla\mathcal{L}(\mathbf{x}_{k})\|>\epsilon$ do
@@ IV Three-Phase LinDistFlow (LinDist3Flow)
where $\mathbb{V}=\begin{bmatrix}v^{a}_{1}&v^{b}_{1}&v^{c}_{1}&\cdots&v^{a}_{n}&v^{b}_{n}&v^{c}_{n}\end{bmatrix}^{\top}$ is the vector of squared voltage magnitudes for each phase $(a,b,c)$, $\mathbb{A}_{3}$ is the network incidence matrix, $\text{bdiag}(\,\cdot\,)$ is the block diagonal operator, $\mathbb{P}=\begin{bmatrix}p^{a}_{1}&p^{b}_{1}&p^{c}_{1}&\cdots&p^{a}_{n}&p^{b}_{n}&p^{c}_{n}\end{bmatrix}^{\top}$ and $\mathbb{Q}=\begin{bmatrix}q^{a}_{1}&q^{b}_{1}&q^{c}_{1}&\cdots&q^{a}_{n}&q^{b}_{n}&q^{c}_{n}\end{bmatrix}^{\top}$ are the active and reactive power injection vectors, and the $\mathbb{H}$ matrices for each line $(i,j)\in\mathcal{E}$ are:
@@
[…]
@@
$\mathbf{g}^{\eta}=\frac{2}{|\mathcal{M}||3\mathcal{N}^{\prime}|}\sum_{m\in\mathcal{M}}\Big(\left.\frac{\partial\mathbb{V}^{OLDF}}{\partial\eta}\right|_{\mathbb{V}^{OLDF}_{m}}\Big)^{\top}\Big(\mathbb{V}_{m}^{OLDF}-\mathbb{V}_{m}^{DF}\Big),$ (12a)